Calculate the total force on charge A. All three charges are 2.5uC. The distance from A to B is 2.0 m and from B to C is 2.0 m and angle B is 90°




1
Expert's answer
2022-03-07T11:45:09-0500

"F_1=\frac{kq^2}{r_1^2}=14.06~mN,"

"F_2=\frac{kq^2}{r_2^2}=7.03~mN,"

"F=\sqrt{F_1^2+F_2^2+2F_1F_2\cos45°}=19.67~mN."